# Exploring simplicity bias in 1D dynamical systems

### v1

@@ Abstract.
Keywords: Dynamical systems; algorithmic probability; simplicity bias; time series; machine learning
@@ 1. Introduction
Predicting the behaviour and trajectories of dynamical systems is a research goal relevant to many fields, including in financial, meteorological, medical, and numerous other applications. Such predictions often rest on a detailed analysis of the underlying model, but recent years have seen machine learning advances provide powerful tools for the prediction and analysis of physical phenomena, including dynamical systems. These machine learning predictions rely on data observations rather than details of the model [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].
An open and important question, then, is the extent to which machine learning methods are applicable in dynamical systems problems, and what kinds of limitations or advantages this relatively novel approach may have. Since information theory and machine learning are inextricably linked and have even been referred to as “two sides of the same coin” [35], one perspective on these questions is to consider whether computation and information processing — fundamental components of machine learning — themselves might have limits of applicability, or in some other way constrain or inform dynamical behaviours.
Some fascinating examples of such limits are known in relation to uncomputability of trajectories, meaning that some properties of dynamical trajectories may not be possible to calculate, even in principle. In a seminal paper, Moore [36] proved that the motion of a single particle in a potential can (in some specific settings) be sufficiently complex as to simulate a Turing machine, and thereby yield trajectories with uncomputable quantitative properties. More recently, Watson et al. [37] proved for an example many-body system that even if the exact initial values of all system parameters were known, the renormalisaton group trajectory and resultant fixed point is impossible to predict. This kind of unpredictability is stronger than the unpredictability of (merely) chaotic dynamics, in which the limiting factor is the accuracy with which initial conditions can be measured. See also Wolfram [38, 39], Svozil [40], Lloyd [41], and Aguirre et al. [42] for more discussion of (un)computability and (un)predictability in physical systems. Naturally, if the dynamics of some system cannot be computed even in principle, machine learning approaches to prediction in these settings will by definition also be severely limited.
As another potential limitation of machine learning methods for the analysis of dynamical systems, also derived from theoretical computer science, consider the work of Cubitt et al. [43] who showed that extracting dynamical equations from data is NP-hard. This means that there is no efficient way to do this in general (assuming that $P\neq NP$, as is widely believed). How significant a limitation this is in practice is debatable, because as they note, the experimental datasets we typically encounter in practice seem to be fairly easy to analyse (while the reason for this is an open question). Regardless, the findings illustrate how fundamental limitations, derived from computer science, may shed light on challenges which face machine learning methods.
These intriguing studies cited above motivate more exploration of the application of theoretical computer science to physics and dynamics [44]. With this broader goal in mind, we consider here another branch of theoretical computer science, algorithmic information theory [45, 46, 47] (AIT), and specifically algorithmic probability, and how arguments inspired by these might facilitate the prediction or at least bounding the behaviour of dynamical systems.
In recent years several studies of simplicity bias have been made in input-output maps, in which a general inverse relationship between the complexity of outputs and their respective probabilities has been observed [48, 49]. More specifically, using arguments based on AIT an upper bound on the probability $P(x)$ of observing output pattern $x$ was presented [48], with the bound depending on the estimated Kolmogorov complexity of $x$. The upper bound implies that complex output patterns must have low probabilities, while high probability outputs must be simple. Example systems where simplicity bias has been observed include RNA structures [48, 50], differential equation solutions [48], finite state transducers [49], time series patterns in natural data [51], natural protein structures [52], among others.
@@
Although not restricted to studying binary strings, most AIT results are framed in the context of binary strings and hence applications are easier in the same framework. Thus in this work we will study simplicity bias in digitised binary trajectories of these example 1D dynamical systems. Our main conclusions are that simplicity bias appears the logistic map, the Gauss (“mouse”) map, and the sine map, and hence we suggest that simplicity bias may also appear in natural dynamical systems more generally.
@@ 2.2. Digitised map trajectories
Figure 2. Bifurcation diagram for the logistic map. In (a), the diagram for parameters $\mu\in$(0, 4.0]; and in (b), for values $\mu\in$(2.9, 4.0]. The value 0.5 has been highlighted in red, to indicate the threshold cut off used to digitise trajectories by a value o 00 if the output is below the threshold, and a value of $1$ if it is greater than or equal to the threshold.
@@ 3.1. Logistic map
The logistic map in Eq. (6) is possibly the best known and most iconic 1D maps studied in dynamical systems and chaos theory. To our knowledge, there have not been many studies of the complexity of digitised trajectories of the logistic map. A notable exception is the work by Kaspar and Schuster [90] who also studied the estimated complexity of resulting binary string trajectories. However their work was fundamentally different in that it was not concerned with simplicity bias or in estimating the probability of different outputs.
@@ 3.1.1. Parameter intervals
In Figure 2(a) the bifurcation diagram of the logistic map is shown, in which for different values of $\mu$ the asymptotic $x_{k}$ values are depicted. The diagram shows fixed points, oscillations, and non-period behaviour. Also added is the value $0.5$ as a red line, highlighting the digitising threshold. It is known [63] that if $\mu\in[0.0,1.0]$, then the trajectories tend to 0. Because we truncate at 0.5, therefore the corresponding binary string would be $x=...0000$. If $\mu\in(1.0,2.0]$ then the trajectories tend to $1-(1/\mu)$, which means that we expect to see binary strings $x=...0000$. For $\mu\in(2.0,3.0]$, the resulting pattern would be or $x=...1111$ because $1-(1/\mu)$ is greater than 0.5. For $\mu$ from 3.0 to $\approx$3.3, we still expect $x=...1111$ because although the first bifurcation appears at $\mu=$3.0, and oscillations with period two begin, until about 3.3 both values of the oscillations are larger than 0.5. Figure 2(b) shows the same bifurcation diagram, but zoomed in on the larger values of $\mu$. From about 3.3 to about 3.5, oscillations between two or four values appear, and larger values of $\mu$ can yield oscillation periods of 8, 16, etc., which will yield patterns such as $x=0101...$. Chaotic trajectories do not occur until [63] $\mu\geq 3.56994567...\approx 3.57$ (https://oeis.org/A098587), and so the $\mu$ interval 3.0 to about 3.5699 contains period-doubling bifurcations, in which oscillations of exponentially increasing frequency occur but no truly chaotic trajectories. Finally, for $\mu$ between $\approx$3.57 and 4.0, more complex and chaotic patterns can emerge, but also ‘islands’ of stability appear in this interval, as can be observed in Figure 2(b).
In our numerical simulations, we will separately investigate various intervals for $\mu$, namely $\mu$ sampled uniformly from (0, 4.0], [3.0, 4.0], [3.57, 4.0], and finally also fixing $\mu=4.0$. The motivation for choosing these intervals is as follows: Given that the interval $\mu\in(0.0,4.0]$ is the standard range in which the map is studied, the most ‘natural’ sampling strategy is to sample uniformly across this interval. As for choosing intervals [3.0, 4.0] and [3.57, 4.0], these are interesting test cases because we can expect to see complex patterns appearing more frequently in these intervals. Finally, when fixing $\mu=4.0$ most trajectories will be highly complex, and so we might expect simplicity bias to disappear.
@@ 3.1.2. Connection of simplicity and probability
Figure 3. Simplicity bias in the digitised logistic map from random samples with $x_{0}\in(0,1)$ and $\mu$ sampled in different intervals. Each blue datapoint corresponds to a different binary digitised trajectory $x$ of length 25 bits. The black line is the upper bound prediction of Eq. (3). (a) Clear simplicity bias for $\mu\in$(0, 4.0] with $P(x)$ closely following the upper bound, except for low frequency and high complexity outputs which suffer from increased sampling noise; (b) simplicity bias is still present for $\mu\in$[3.0, 4.0]; (c) the distribution of $P(x)$ becomes more flat (less biased) and simplicity bias is much less clear when $\mu\in$[3.57, 4.0] due to constraining the sampling to $\mu$-regions more likely to show chaos; (d) the distribution of $P(x)$ is roughly uniform when using $\mu=4.0$, with almost no bias, and hence no possibility of simplicity bias.
@@ 3.1.3. Simplicity bias appears when bias appears
Figure 3(a) shows that the upper bound prediction (black line) agrees remarkably well with the probability and complexity data for different binary string output (blue dots); we see that the logistic map displays simplicity bias when uniform sampling $\mu\in(0.0,4.0]$, such that high probability outputs have low complexities, and complex outputs have low probability. The gradient prediction ($a=1$) of the black line used no information about the map, except that we assumed that almost all $2^{n}$ outputs of length $n$ bits are realisable. Note that many output strings fall below the upper bound prediction as we expected from earlier studies [48, 86], but nonetheless it is known that randomly generated outputs tend to be close to the bound [49]. Put differently, even though many output strings (blue dots) appear to be far below the bound, most of the probability mass for each complexity value is concentrated close to the bound. Thus, this simple bound predicts $P(x)$ quite accurately by using the complexity values of output strings yet while otherwise completely ignoring the details of the underlying dynamical system.
In Figure 3(b) we make a similar plot to panel (a) except that we restrict the sampling to $\mu\in[3.0,4.0]$. The qualitative pattern in (b) is similar to (a), although simplicity bias is slightly less clear than in (a). For Figure 3(c) the sampling interval was chosen so that $\mu\in[3.57,4.0]$ which is the region containing some truly chaotic trajectories [89]. Both bias and simplicity bias are less clear in this plot. In Figure 3(d) a plot is given in which $\mu$ is no longer sampled, but is fixed at $\mu=4.0$, so that almost all $x_{0}$ values yield chaotic trajectories. As expected, there is very little bias in the distribution, i.e. the distribution of $P(x)$ is roughly uniform, and hence no simplicity bias can be observed. Figure 3(d) still shows some very simple strings, which can be understood from the fact that $x_{i}\approx 0$ $\Rightarrow x_{i+j}\approx 0$ for $j\approx 1$ even if the trajectories may become much more complex for $j\gg 1$. In other words, if the trajectory reaches a value close to zero, then it tends to stay close to zero for several iterations. These initial short trajectories with many 0s will have very low complexity.
@@
See the Appendix A for an illustration of the effects of using the different values of $n$.
Figure 4. The distribution $P(\tilde{K}(x)=r)$ of output complexity values, with $x_{0}\in(0,1)$ and $\mu$ sampled from different intervals. (a) A roughly uniform complexity distribution for $\mu\in$(0.0, 4.0], with some bias towards lower complexities; (b) Close to uniform distribution of complexities for $\mu\in$[3.0, 4.0]; (c) the distribution leans to higher complexities when $\mu\in$[3.57, 4.0]; (d) the distribution is biased to higher complexities values when $\mu=4.0$; (e) for comparison, purely random binary strings of length 25 bits were generated. The distributions of complexity values in (d) and (e) are very similar, but (a-c) show distinct differences. Calculating and comparing $P(K)$ is an efficient way of checking how simplicity biased a map is.
@@ 3.1.4. Distribution of complexities
To investigate this distribution in the logistic map, in Figure 4 we plot the distribution of complexities for the different sampling intervals of $\mu$. As can be seen in Figure 4(a), when sampling from [0.0, 4.0] there is a bias towards lower complexities. As for (b), when $\mu$ is sampled from [3.0, 4.0] the distribution of complexities is roughly uniform (at least on a log scale). In Figure 4(c), the distribution is also somewhat uniform but peaks around medium complexity values. In (d), there is a bias towards higher complexity values. For comparison, in (e) we also plot the distribution of complexities resulting from sampling purely random binary strings, and this distribution is very similar to that in (d) obtained when $\mu=4.0$. It is noteworthy that in some of these cases, while there is some evidence of bias toward higher or lower complexities, the distributions still display spread and are not narrowly peaked (at least on a log scale). We note that if the logistic map produced binary output strings with a uniform probability over strings, in theory the frequency would exponentially with complexity $r$, however, for Lempel-Ziv complexity with short strings, the distribution is not quite exponential, see [48, 57].
@@ 3.1.6. Pre-chaotic regime
Figure 5. Simplicity bias in (a) the logistic map with $\mu$ sampled in [0.0, 3.5699], which is the non-chaotic period doubling regime; (b) the Gauss map; and (c) the sine map.
@@ 4. Discussion
Arguments inspired by algorithmic information theory (AIT) predict that in many input-output maps, there will be probabilistic bias with complex patterns having exponentially low probabilities, and some simple patterns having high probability; this phenomenon is known as simplicity bias [48]. Here, we numerically investigated the presence of simplicity bias in digitised trajectories arising from iterations of the logistic map, Gauss map, sine map, Bernoulli map, and tent map. By digitising the real-valued trajectories, we studied the probability and complexity of the resulting binary strings. Our main conclusions are that (i) we observe simplicity bias in the logistic map, Gauss map, and sine map, and also that in some cases the probability of resulting binary strings can be predicted a priori with surprising accuracy; and (ii) we do not observe simplicity bias in the trajectories of the Bernoulli map and tent map, nor indeed any bias at all.
@@
Due to the qualitatively different behaviours exhibited by the logistic map for different $\mu$ values, we separately studied different regimes by sampling $\mu$ from (0.0, 4.0], (3.0, 4.0], (3.57, 4.0] and also $\mu=4.0$. In general, simplicity bias and upper bound prediction accuracy was higher for $\mu$ sampled across the full range (0.0, 4.0] and decreased for smaller ranges, until completely disappearing for $\mu=4.0$. The logistic map is perhaps the most iconic example of a dynamical system in chaos theory, and has been very extensively studied for decades. Here we report a novel finding relating to this map, and one that is not (merely) a subtle higher-order effect, but rather a strong effect related to order-of-magnitude variations in pattern probability. This finding is also interesting given that we did not necessarily expect to observe simplicity bias when outputs can be pseudo-random (Cf. ref. [48]). It appears that in this map, pseudo-random outputs are sufficiently rare that they do not cause strong violations of the simplicity bias upper bound. Additionally, we found simplicity bias can be ‘tuned’ via altering the $\mu$ interval: sampling from the full interval (0.0, 4.0] yields a biased (low entropy) distribution over output strings along with simplicity bias, while sampling from higher values of $\mu\approx 4.0$ yields low bias (high entropy) distributions and little or no simplicity bias. While we observe simplicity bias similar to that predicted by AIT arguments in some of these maps, we want to clarify that we do not think that we have in any way proven that these patterns are in fact directly linked to the AIT arguments. For that much more work is needed. Nevertheless, we argue that studying probability-complexity relationships, and looking for patterns such as simplicity bias may be a fruitful perspective for dynamical systems research.
The motivations for this work were to explore new examples of simplicity bias in physics and specifically dynamical systems, test the boundaries of which types of maps show simplicity bias, and explore how information theory and algorithmic probability can inform dynamical systems research. By extension, this work expands on the project of investigating the interaction between machine learning and dynamical systems because machine learning is intimately connected to information theory and algorithmics. Indeed, it may be that dynamical systems that are easy to learn, are precisely those that exhibit some kind of simplicity bias which matches the inbuilt Occam’s razor like inductive bias of DNNs [57].
@@
Mathematicians and physicists are fascinated by studying simple systems which can exhibit complex patterns, even while they are not technically random. In this context, Wolfram [39] investigated simple automata, and showed that some have high levels of computing power, leading to his conjecture that many natural systems can be Turing complete, i.e., that they can implement arbitrary algorithms, and hence produce complex patterns. Despite this focus on complexity, we observe in our work here that complexity is in a sense not actually that common: Even though the logistic map is famous precisely due to its ability to produce complex behaviour, within the space of possible parameters, i.e., $\mu,x_{0}\in\mathcal{R}$ and even restricting to $\mu\in(0.0,4.0]$ and $x_{0}\in(0,1)$, chaos and ‘complexity’ only occur rarely (assuming uniform sampling of the parameter space). This observation accords with the fact that while Wolfram highlighted some rule sets that produce randomness, most of his automata rules sets do not produce complex pseudo-random patterns [39]. Coe et al.  [103] analytically studied the question of when automata produce ‘random’ and complex dynamics, and also found that relatively few are random.
@@
The study of random dynamical systems — in which dynamical systems are randomly perturbed in some way — is quite well established [104], including specifically dynamical systems with additive noise [105]. While deterministic dynamical systems have been studied due to their relevance to modelling problems in e.g., ecology, physics, and economics, randomly perturbed dynamics are also common in science and hence important to study, and arise naturally in modelling physical systems (e.g., [106]). This motivates studying simplicity bias also in such random dynamical systems in future work. As a separate motivation, in the deterministic logistic map, many sampled $\mu$ values yield trajectories which quickly converge to fixed points or other fairly trivial patterns. However, by introducing some small random noise these trajectories may be prevented from converging into trivial patterns, and therefore may show simplicity bias even while the deterministic counterpart does not. One potential map to utilise would be the logistic map with small additive noise, as studied by Sato et al. [107].
@@
In this work we have used 1D maps including the logistic map as toy systems to explore simplicity bias in dynamical systems. While our work has been an empirical investigation, the connection between AIT and dynamical systems has also received some earlier limited attention from an analytical perspective (e.g., [108, 109, 110]). In future work it may be fruitful to investigate which, if any, properties of the logistic (or other chaotic) map can be predicted or bounded using the simplicity bias bound. That is, if we assume that the simplicity bias bound holds, what other dynamical properties of the map might follow from this? Even if these properties are already known, it would still be insightful to see if they could be derived (perhaps more simply) from AIT arguments. Further, the presence of simplicity bias in some of the 1D map may lead to using the simplicity bias bound as a trajectory prediction method. This is related to the a priori prediction of natural time series patterns attempted by Dingle et al. [51]. More generally, exploring the use of arguments inspired by AIT in dynamical systems and chaos research is an attractive and potentially fruitful research avenue.

### v2

@@ Abstract.
Keywords: Dynamical systems; algorithmic probability; simplicity bias; time series
@@ 1. Introduction
In recent years several studies of simplicity bias have been made in input-output maps, in which a general inverse relationship between the complexity of outputs and their respective probabilities has been observed [1, 2]. More specifically, using arguments inspired by algorithmic information theory [3, 4, 5] (AIT), and specifically algorithmic probability, an upper bound on the probability $P(x)$ of observing output pattern $x$ was presented [1], with the bound depending on the estimated Kolmogorov complexity of $x$. The upper bound implies that complex output patterns must have low probabilities, while high probability outputs must be simple. Example systems where simplicity bias has been observed include RNA structures [1, 6], differential equation solutions [1], finite state transducers [2], time series patterns in natural data [7], natural protein structures [8], among others.
@@
Although not restricted to studying binary strings, most AIT results are framed in the context of binary strings and hence applications are easier in the same framework. Thus in this work we will study simplicity bias in digitised binary trajectories of these example 1D dynamical systems. Our main conclusions are that simplicity bias appears in the logistic map, the Gauss (“mouse”) map, and the sine map, and hence we suggest that simplicity bias may also appear in natural dynamical systems more generally. As a broader context of motivation, this work contributes to research is the intersection of dynamical systems, and AIT, and machine learning.
@@ 2.2. Digitised map trajectories
Figure 2. Bifurcation diagram for the logistic map. In (a), the diagram for parameters $\mu\in$(0, 4.0]; and in (b), for values $\mu\in$(2.9, 4.0]. The value 0.5 has been highlighted in red, to indicate the cut off threshold used to digitise trajectories by a value of 00 if the output is below the threshold, and a value of $1$ if it is greater than or equal to the threshold.
@@ 3.1. Logistic map
The logistic map in Eq. (6) is possibly the best known and most iconic 1D map studied in dynamical systems and chaos theory. To our knowledge, there have not been many studies of the complexity of digitised trajectories of the logistic map. A notable exception is the work by Kaspar and Schuster [41] who also studied the estimated complexity of resulting binary string trajectories. However their work was fundamentally different in that it was not concerned with simplicity bias or in estimating the probability of different outputs.
@@ 3.1.1. Parameter intervals
In Figure 2(a) the bifurcation diagram of the logistic map is shown, in which for different values of $\mu$ the asymptotic $x_{k}$ values are depicted. The diagram shows fixed points, oscillations, and non-period behaviour. Also added is the value $0.5$ as a red line, highlighting the digitising threshold. It is known [14] that if $\mu\in[0.0,1.0]$, then the trajectories tend to 0. Because we truncate at 0.5, therefore the corresponding binary string would be $x=...0000$. If $\mu\in(1.0,2.0]$ then the trajectories tend to $1-(1/\mu)$, which means that we expect to see binary strings $x=...0000$. For $\mu\in(2.0,3.0]$, the resulting pattern would be $x=...1111$ because $1-(1/\mu)$ is greater than 0.5. For $\mu$ from 3.0 to $\approx$3.3, we still expect $x=...1111$ because although the first bifurcation appears at $\mu=$3.0, and oscillations with period two begin, until about 3.3 both values of the oscillations are larger than 0.5. Figure 2(b) shows the same bifurcation diagram, but zoomed in on the larger values of $\mu$. From about 3.3 to about 3.5, oscillations between two or four values appear, and larger values of $\mu$ can yield oscillation periods of 8, 16, etc., which will yield patterns such as $x=0101...$. Chaotic trajectories do not occur until [14] $\mu\geq 3.56994567...\approx 3.57$ (https://oeis.org/A098587), and so the $\mu$ interval 3.0 to about 3.5699 contains period-doubling bifurcations, in which oscillations of exponentially increasing frequency occur but no truly chaotic trajectories. Finally, for $\mu$ between $\approx$3.57 and 4.0, more complex and chaotic patterns can emerge, but also ‘islands’ of stability appear in this interval, as can be observed in Figure 2(b).
In our numerical simulations, we will separately investigate various intervals for $\mu$, namely $\mu$ sampled uniformly from (0.0, 4.0], [3.0, 4.0], [3.57, 4.0], and finally also fixing $\mu=4.0$. The motivation for choosing these intervals is as follows: Given that the interval $\mu\in(0.0,4.0]$ is the standard range in which the map is studied, the most ‘natural’ sampling strategy is to sample uniformly across this interval. As for choosing intervals [3.0, 4.0] and [3.57, 4.0], these are interesting test cases because we can expect to see complex patterns appearing more frequently in these intervals. Finally, when fixing $\mu=4.0$ most trajectories will be highly complex, and so we might expect simplicity bias to disappear.
@@ 3.1.2. Connection of simplicity and probability
Figure 3. Simplicity bias in the digitised logistic map from random samples with $x_{0}\in(0,1)$ and $\mu$ sampled in different intervals. Each blue datapoint corresponds to a different binary digitised trajectory $x$ of length 25 bits. The black line is the upper bound prediction of Eq. (3). (a) Clear simplicity bias for $\mu\in$(0.0, 4.0] with $P(x)$ closely following the upper bound, except for low frequency and high complexity outputs which suffer from increased sampling noise; (b) simplicity bias is still present for $\mu\in$[3.0, 4.0]; (c) the distribution of $P(x)$ becomes more flat (less biased) and simplicity bias is much less clear when $\mu\in$[3.57, 4.0] due to constraining the sampling to $\mu$-regions more likely to show chaos; (d) the distribution of $P(x)$ is roughly uniform when using $\mu=4.0$, with almost no bias, and hence no possibility of simplicity bias.
@@ 3.1.3. Simplicity bias appears when bias appears
Figure 3(a) shows that the upper bound prediction (black line) agrees remarkably well with the probability and complexity data for different binary string output (blue dots); we see that the logistic map displays simplicity bias when uniform sampling $\mu\in(0.0,4.0]$, such that high probability outputs have low complexities, and complex outputs have low probability. The gradient prediction ($a=1$) of the black line used no information about the map, except that we assumed that almost all $2^{n}$ outputs of length $n$ bits are realisable. An upper-bound fit to the $\log_{10}P(x)$ data gives the slope as -0.18. Note that many output strings fall below the upper bound prediction as we expected from earlier studies [1, 37], but nonetheless it is known that randomly generated outputs tend to be close to the bound [2]. Put differently, even though many output strings (blue dots) appear to be far below the bound, most of the probability mass for each complexity value is concentrated close to the bound. Thus, this simple bound predicts $P(x)$ quite accurately by using the complexity values of output strings yet while otherwise completely ignoring the details of the underlying dynamical system.
In Figure 3(b) we make a similar plot to panel (a) except that we restrict the sampling to $\mu\in[3.0,4.0]$. The qualitative pattern in (b) is similar to (a), although simplicity bias is slightly less clear than in (a). An upper-bound fit to the $\log_{10}P(x)$ data gives the slope as -0.17. For Figure 3(c) the sampling interval was chosen so that $\mu\in[3.57,4.0]$ which is the region containing some truly chaotic trajectories [40]. Both bias and simplicity bias are less clear in this plot. An upper-bound fit to the $\log_{10}P(x)$ data gives the slope as -0.04. In Figure 3(d) a plot is given in which $\mu$ is no longer sampled, but is fixed at $\mu=4.0$, so that almost all $x_{0}$ values yield chaotic trajectories. As expected, there is very little bias in the distribution, i.e., the distribution of $P(x)$ is roughly uniform, and hence no simplicity bias can be observed. An upper-bound fit to the $\log_{10}P(x)$ data gives the slope as 0.02. Figure 3(d) still shows some very simple strings, which can be understood from the fact that $x_{i}\approx 0$ $\Rightarrow x_{i+j}\approx 0$ for $j\approx 1$ even if the trajectories may become much more complex for $j\gg 1$. In other words, if the trajectory reaches a value close to zero, then it tends to stay close to zero for several iterations. These initial short trajectories with many 0s will have very low complexity.
@@
See the Appendix A for an illustration of the effects of using the different values of $n$. See Appendix B for the same plots as in Figure 3, but with semi-transparent data points to highlight the distribution of data points.
Figure 4. The distribution $P(\tilde{K}(x)=r)$ of output complexity values, with $x_{0}\in(0.0,1.0)$ and $\mu$ sampled from different intervals. (a) A roughly uniform complexity distribution for $\mu\in$(0.0, 4.0], with some bias towards lower complexities (mean is 3.4 bits); (b) Close to uniform distribution of complexities for $\mu\in$[3.0, 4.0], mean is 10.3 bits; (c) the distribution leans to higher complexities when $\mu\in$[3.57, 4.0], mean is 14.1 bits; (d) the distribution is biased to higher complexities values when $\mu=4.0$ (mean is 16.4 bits); (e) for comparison, purely random binary strings of length 25 bits were generated (mean is 16.2 bits). The distributions of complexity values in (d) and (e) are very similar, but (a-c) show distinct differences. Calculating and comparing $P(K)$ is an efficient way of checking how simplicity biased a map is.
@@ 3.1.4. Distribution of complexities
To investigate this distribution in the logistic map, in Figure 4 we plot the distribution of complexities for the different sampling intervals of $\mu$. As can be seen in Figure 4(a), when sampling from [0.0, 4.0] there is a bias towards lower complexities. As for (b), when $\mu$ is sampled from [3.0, 4.0] the distribution of complexities is roughly uniform (at least on a log scale). In Figure 4(c), the distribution is also somewhat uniform but peaks around medium complexity values. In (d), there is a bias towards higher complexity values. For comparison, in (e) we also plot the distribution of complexities resulting from sampling purely random binary strings, and this distribution is very similar to that in (d) obtained when $\mu=4.0$. It is noteworthy that in some of these cases, while there is some evidence of bias toward higher or lower complexities, the distributions still display spread and are not narrowly peaked (at least on a log scale). We note that if the logistic map produced binary output strings with a uniform probability over strings, in theory the frequency would grow exponentially with complexity $r$, however, for Lempel-Ziv complexity with short strings, the distribution is not quite exponential, see [1, 42].
@@ 3.1.6. Pre-chaotic regime
Figure 5. Simplicity bias in (a) the logistic map with $\mu$ sampled in [0.0, 3.5699], which is the non-chaotic period doubling regime (upper bound fitted slope is -0.17); (b) the Gauss map (upper bound fitted slope is -0.13); and (c) the sine map (upper bound fitted slope is -0.17).
@@ 4. Discussion
Arguments inspired by algorithmic information theory (AIT) predict that in many input-output maps, strongly non-uniform probability distributions over outputs will result, with complex patterns having exponentially low probabilities, and some simple patterns having high probability; this phenomenon is known as simplicity bias [1]. Here, we numerically investigated the presence of simplicity bias in digitised trajectories arising from iterations of the logistic map, Gauss map, sine map, Bernoulli map, and tent map. By digitising the real-valued trajectories, we studied the probability and complexity of the resulting binary strings. Our main conclusions are that (i) we observe simplicity bias in the logistic map, Gauss map, and sine map, and also that in some cases the probability of resulting binary strings can be predicted a priori with surprising accuracy; and (ii) we do not observe simplicity bias in the trajectories of the Bernoulli map and tent map, nor indeed any bias at all.
@@
Due to the qualitatively different behaviours exhibited by the logistic map for different $\mu$ values, we separately studied different regimes by sampling $\mu$ from (0.0, 4.0], (3.0, 4.0], (3.57, 4.0] and also $\mu=4.0$. In general, simplicity bias and upper bound prediction accuracy was higher for $\mu$ sampled across the full range (0.0, 4.0] and decreased for smaller ranges, until completely disappearing for $\mu=4.0$. The logistic map is perhaps the most iconic example of a dynamical system in chaos theory, and has been very extensively studied for decades. Here we report a novel finding relating to this map, and one that is not (merely) a subtle higher-order effect, but rather a strong effect related to order-of-magnitude variations in pattern probability. This finding is also interesting given that we did not necessarily expect to observe simplicity bias when outputs can be pseudo-random (Cf. ref. [1]). It appears that in this map, pseudo-random outputs are sufficiently rare that they do not cause strong violations of the simplicity bias upper bound. Additionally, we found simplicity bias can be ‘tuned’ via altering the $\mu$ interval: sampling from the full interval (0.0, 4.0] yields a biased (low entropy) distribution over output strings along with simplicity bias, while sampling from higher values of $\mu\approx 4.0$ yields low bias (high entropy) distributions and little or no simplicity bias. While we observe simplicity bias similar to that predicted by AIT arguments in some of these maps, we want to clarify that we do not think that we have in any way proven that these patterns are in fact directly linked to the AIT arguments. For that, much more work is needed. Nevertheless, we argue that studying probability-complexity relationships, and looking for patterns such as simplicity bias may be a fruitful perspective for dynamical systems research.
@@
The motivations for this work were to explore new examples of simplicity bias in physics and specifically dynamical systems, test the boundaries of relevance of simplicity bias, and explore how information theory and algorithmic probability can inform dynamical systems research. By extension, this work expands on the project of investigating the interaction between machine learning and dynamical systems because machine learning is intimately connected to information theory and algorithmics. The broader context of our work is a research project into the interface of dynamical systems, machine learning, and AIT. Given the strong interest in applying machine learning to the analysis of dynamical systems, an open and important question, then, is the extent to which machine learning methods are applicable in dynamical systems problems, and what kinds of limitations or advantages this relatively novel approach may have. Since information theory and machine learning are inextricably linked and have even been referred to as “two sides of the same coin” [51], one perspective on these questions is to consider whether computation and information processing — fundamental components of machine learning — themselves might have limits of applicability, or in some other way constrain or inform dynamical behaviours. Some fascinating examples of such limits are known in relation to uncomputability of trajectories, meaning that some properties of dynamical trajectories may not be possible to calculate, even in principle. In a seminal paper, Moore [52] proved that the motion of a single particle in a potential can (in some specific settings) be sufficiently complex as to simulate a Turing machine, and thereby yield trajectories with uncomputable quantitative properties. More recently, Watson et al. [53] proved for an example many-body system that even if the exact initial values of all system parameters were known, the renormalisaton group trajectory and resultant fixed point is impossible to predict. This kind of unpredictability is stronger than the unpredictability of (merely) chaotic dynamics, in which the limiting factor is the accuracy with which initial conditions can be measured. See also Wolfram [54, 55], Svozil [56], Lloyd [57], and Aguirre et al. [58] for more discussion of (un)computability and (un)predictability in physical systems. Naturally, if the dynamics of some system cannot be computed even in principle, the accuracy of machine learning approaches to prediction in these settings will be restricted (but see [59]).
@@
Mathematicians and physicists are fascinated by studying simple systems which can exhibit complex patterns, even while they are not technically random. In this context, Wolfram [55] investigated simple automata, and showed that some have high levels of computing power, leading to his conjecture that many natural systems can be Turing complete, i.e., that they can implement arbitrary algorithms, and hence produce complex patterns. Despite this focus on complexity, we observe in our work here that complexity is in a sense not actually that common: Even though the logistic map is famous precisely due to its ability to produce complex behaviour, within the space of possible parameters, i.e., $\mu,x_{0}\in\mathcal{R}$ and even restricting to $\mu\in(0.0,4.0]$ and $x_{0}\in(0.0,1.0)$, chaos and ‘complexity’ only occur rarely (assuming uniform sampling of the parameter space). This observation accords with the fact that while Wolfram highlighted some rule sets that produce randomness, most of his automata rules sets do not produce complex pseudo-random patterns [55]. Coe et al.  [68] analytically studied the question of when automata produce ‘random’ and complex dynamics, and also found that relatively few are random.
@@
The study of random dynamical systems — in which dynamical systems are randomly perturbed in some way — is quite well established [69], including specifically dynamical systems with additive noise [70]. While deterministic dynamical systems have been studied due to their relevance to modelling problems in e.g., ecology, physics, and economics, randomly perturbed dynamics are also common in science and hence important to study, and arise naturally in modelling physical systems (e.g., [71]). This motivates studying simplicity bias also in such random dynamical systems in future work. As a separate motivation, in the deterministic logistic map, many sampled $\mu$ values yield trajectories which quickly converge to fixed points or other fairly trivial patterns. However, by introducing some small random noise these trajectories may be prevented from converging into trivial patterns, and therefore may show simplicity bias even while the deterministic counterpart does not. The relation of simplicity bias to random dynamical systems has been initially studied recently [72], but many questions remain open.
@@
In this work we have used 1D maps including the logistic map as toy systems to explore simplicity bias in dynamical systems. The connection between AIT and dynamical systems has received some earlier attention from an analytical perspective [73, 74, 75], and computational perspective [76, 72]. In future work it may be fruitful to investigate which, if any, properties of the logistic (or other chaotic) map can be predicted or bounded using the simplicity bias bound. That is, if we assume that the simplicity bias bound holds, what other dynamical properties of the map might follow from this? Even if these properties are already known, it would still be insightful to see if they could be derived (perhaps more simply) from AIT arguments. Further, the presence of simplicity bias in some of the 1D map may lead to using the simplicity bias bound as a trajectory prediction method. This is related to the a priori prediction of natural time series patterns attempted by Dingle et al. [7]. Another angle would be to study non-digitised trajectories of dynamical systems, which would require different complexities measures amenable to continuous curves, such as proposed in ref. [77]. More generally, exploring the use of arguments inspired by AIT in dynamical systems and chaos research is an attractive and potentially fruitful research avenue.